Deposit of sulfur in volcanic regions are thought to have resulted from the reaction of H2S and SO2. both are constituents of volcanic gases. The reaction is

2 H2S +SO2 2H2O + 3S

 

Gases in a mixture containing 13.6 g H2S and 28.83 g SO2 are allowed to react.

Which is the limiting reactant?

How many grams of sulfur would be formed based on the amount of the limiting reactant in this mixture?

How many grams of the excess reactant will remain?


1
Expert's answer
2021-11-08T08:24:04-0500

2H2S+SO2--->>2H2O + 3S.

Molar mass of SO2= 64.066

28.83/64.066= 0.45

Molar mass of H2S= 34.1 g/mol

13.6/34.1=0.399mol

Molar mass of S= 32.065

3×0.399/2= 0.599

0.45×3=1.35

1.35+0.599=1.949

= 1.949×32.065=62.495g